# Topic-Specific Classifiers are Better Relevance Judges than Prompted LLMs

### v1

@@ Abstract
The unjudged document problem, where pooled test collections have incomplete relevance judgments for evaluating new retrieval systems, is a key obstacle to the reusability of test collections in information retrieval. While the de facto standard to deal with the problem is to treat unjudged documents as non-relevant, many alternatives have been proposed, including the use of large language models (LLMs) as a relevance judge (LLM-as-a-judge). However, this has been criticized as circular, since the same LLM can be used as a judge and as a ranker at the same time. We propose to train topic-specific relevance classifiers instead: By finetuning monoT5 with independent LoRA weight adaptation on the judgments of a single assessor for a single topic’s pool, we align it to that assessor’s notion of relevance for the topic. The system rankings obtained through our classifier’s relevance judgments achieve a Spearmans’ $\rho$ correlation of >0.95>0.95 with ground truth system rankings. As little as 128 initial human judgments per topic suffice to improve the comparability of models, compared to treating unjudged documents as non-relevant, while achieving more reliability than existing LLM-as-a-judge approaches. Topic-specific relevance classifiers thus are a lightweight and straightforward way to tackle the unjudged document problem, while maintaining human judgments as the gold standard for retrieval evaluation. Code, models, and data are made openly available. [^1] https://github.com/lgienapp/ecir26-topic-adapters-as-judges
1 Introduction
@@
A retrieval experiment usually consists of a large collection of documents, several dozen topics, and manual relevance judgments of a sample of the documents [4, 30]. For large collections, the manual effort involved usually prevents a complete assessment of all documents for all topics. Furthermore, most documents in a collection are typically not relevant to a specific topic. To find a compromise between evaluation costs and experimental validity, it is therefore necessary to build a focused subset of the documents for each topic, obtained by combining the top-$k$ documents from the various search engines to be evaluated, where $k$ can range from a handful to a few hundred documents per search engine.
@@
While this evaluation setup allows for a valid and fair comparison between the systems that initially contribute to the document pool of each topic, any new systems that are implemented after the experiment is complete are at a disadvantage [16]. These often retrieve documents that are not yet included in any given topic’s pool, leading to the so-called unjudged document problem, namely the still mostly unresolved question of how to deal with such documents. Heuristic solutions such as simply omitting these documents, or deeming them non-relevant, have been proposed and frequently applied until today. However, none of the heuristics put new and initial systems on an equal footing. As research progresses and more unjudged but potentially relevant documents are uncovered by new retrieval systems, past evaluation experiments and test collections lose the capability to reliably distinguish between them, diminishing their reusability. At the same time, updating a test collection with new judgments is costly and the new judgments for a topic may be incongruent with those of the previous judges [31].
@@
Recent work has proposed to simply prompt instruction-tuned large language models (LLMs) to supplement or completely replace human judgments [27, 2, 19, 1, 28], collectively called LLM-as-a-judge approaches. However, any LLM used as a judge for evaluation (together with its judgment prompt) can simultaneously be used as a ranker in a retrieval system [9]. This may lead to an LLM confirming its own ranking decisions during relevance judgment, biasing evaluation results in favor of a particular LLM. Moreover, since LLMs inevitably encode social biases picked up from their pre-training data, evaluating retrieval models with LLMs may amplify them in retrieval models [8].
@@
We show that the key to address both, the unjudged document problem, and the bias due to using LLMs for both ranking and evaluation, is to train one relevance classifier for each topic based on the judgments of a single human assessor. A topic-specific classifier can only be used to judge documents for that topic, and it ideally closely approximates the notion of relevance originally applied by the human assessor. To demonstrate the effectiveness of our approach, we conduct experiments using the TREC Robust04 [29], the TREC Deep Learning 2019 [7], and the TREC Deep Learning 2020 [6] passage ranking test collections. We show that the system rankings produced by our topic-specific relevance classifiers strongly correlates with the rankings produced by the available human judgments. We analyze our approach with respect to shallow vs. deep human judgments, and in comparison to zero and few-shot LLM-as-a-judge approaches. Our relevance classifier correlates reliably with human judgments even under limited judgment depth, while outperforming LLM-as-a-judge approaches. These results strongly indicate the feasibility of using topic-specific relevance classifiers to automate retrieval evaluation after bootstrapping one classifier for every topic with a sufficient amount of gold standard human judgments.
2 Related Work
@@
We review the literature, relating our work to attempts at automating relevance judgments from early to modern information retrieval, with a particular focus on recent approaches, and how they differ from ours.
@@
The notion of treating document ranking as a classification problem has been established early in information retrieval with the binary independence model [22]. Despite the fact that this paradigm to ranking was, for a long time, less effective than other approaches, the idea of exploiting relevance judgments to tune relevance classifiers has also been proposed a long time ago, using logistic regression [10] and investigating active learning as a means to minimize the number of human judgments [14]. Unlike our approach, this notion of ranking attempts to train a generic relevance classifier for ranking instead of judging, which requires a lot of judgments, since the notion of relevance may differ wildly between topics and between human assessors. Later learning to rank approaches succeeded only once large query logs have been identified as source of implicit relevance feedback [13].
Topic-specific relevance classifiers for ranking have been studied at the TREC Filtering tracks 1994 to 2002 [23], where participants trained many different kinds of classifiers to filter (classify) documents by their relevance with respect to various different settings, ranging from decision tress to SVMs to small neural networks. And since at TREC, one assessor typically is assigned one topic for judgment, this implicitly means that all these classifiers also approximated the notion of relevance of that specific assessor, just like we propose in this paper.
Active Learning of Relevance Judgments
Automating the collection of relevance judgments has been a long-standing problem in information retrieval [26]. It has been investigated to prioritize documents for judgment and to determine when to stop the assessment [24], to enable in eDiscovery through technology-assisted review [5], at the TREC Total Recall tracks 2015 and 2016 [11], and to train classifiers to judge documents outside domain-specific tasks [32]. However, using active learning typically requires many rounds of human feedback to learn to reliably identify relevant documents, and our goal is not to reduce the number of judgments needed below what is already being done at TREC today.
@@
With the rapid rise of LLMs, they have been demonstrated to have a general sense of relevance [18, 17]. Many studies have thus investigated their applicability at collecting relevance judgments [27, 21, 19, 1] or even fully generating synthetic test collections [20]. The general technique is to develop a prompt that encodes assessment instructions and to input this prompt along a query and unjudged documents. We argue that instead, topic specialization (i.e., learning from as much topic expertise as can be obtained) is key to automating relevance judgment in Cranfield-style evaluations, which the current zero-shot approach of LLM-as-a-judge forgoes in favor of generic instructions. While these studies have shown LLM judges to be correlated with human assessment, there are three main issues that current approaches do not address, namely that an LLM-as-a-judge (1) exhibits biases and reliability issues [25], (2) is highly susceptible to prompt variations [2], and (3) can lead to invalid conclusions about evaluation, especially due to circularity [9].
@@
The idea to predict relevance judgments and expand the overall set of relevance judgments was first proposed by Büttcher et al. [3]. The general idea was to train a binary SVM classifier using TF–IDF features from documents to determine if a document is relevant or not. There are three main issues with these experiments that we address in this paper: (1) Their approach did not work well for deep assessments. Our approach effectively predicts relevance at much deeper depth. (2) Their approach relies heavily on lexical features. Today, effective ranking models use much more complex, semantic scoring functions and our approach models semantics much more effectively. (3) Their approach trains classifiers from scratch, while we opt for adaptation of existing ranking models, leveraging their strong relevance priors.
3 Topic-Specific Relevance Classifiers for Judging
@@
We leverage pretrained ranking models as initialization rather than training classifiers from scratch. Models pretrained on large-scale relevance data provide strong semantic representations and implicit priors over document-query relevance [18, 17]. We experiment with base ranking models from three archetypes: bi-encoders (representation-based), cross-encoders (interaction-based), and mono-decoders (generative scoring). Yet, full fine-tuning per topic is computationally prohibitive and produces large model weights unfit for distribution. We thus rely on LoRA [12] to inject topic-specific behavior while keeping base model weights frozen. LoRA learns a low-rank approximation of the weight modifications attained during training. This allows each topic to receive independent adapter weights with no parameter sharing, preventing cross-topic interference. It further allows for lightweight distribution: one base model, many shareable topic-specific adapters.
4 Experimental Setup
We conduct three experiments. The first investigates if, given a deep pool of judged data, training judge models is feasible; the second investigates how this changes if the pool changes to a shallow one, and how the predicted relevance judgments produced by a finetuned LoRA adapter (‘judge adapter’) impact ranking evaluation results. Finally, the third investigates how the approach compares to LLM-as-a-judge methods, and if those can likewise benefit from topic-specific alignment through few-shot prompting.
@@ 4 Experimental Setup
We select 47 topics with at least 100 relevant documents from the deeply judged official eval set of the TREC Robust04 dataset [29]. Annotated graded relevance labels are cast to binary relevance. This yields, on average, $1350.31(444.23)$ non-relevant and $183.02(78.76)$ relevant documents per topic. We divide the available judgments in each topic into a train and a test set with a 80/20 split, with stratified sampling per label to ensure similar relevancy distribution in both sets. Then, for each of the three pointwise model archetypes bi-encoder, cross-encoder, and mono-decoder, we select representative pretrained models as listed in Table˜1. We separately finetune topic-specific judge models for each topic and archetype, over 10 epochs, with a batch size of 64 query/doc pairs of a combined sequence length of 512 tokens, and a learning rate of 1e-4. We train two version per topic: a native finetuning of the full model weights, and a LoRA adapter [12] with parameters $r=64$ and $\alpha=128$. The number of trainable parameters in each setting and for each model is also listed in Table˜1. Given the high class imbalance in each topic, we use a weighted MSE loss, with weights 0.95 and 0.05 for relevant/non-relevant labels. This corresponds to the inverse average relevance distribution in judgments.
Table 1: Overview of the models investigated as topic-specific judges in this paper.
@@
Bi-Encoder hf.co/sentence-transformers/all-MiniLM-L6-v2 22.7 M 5.4 M
Cross-Encoder hf.co/cross-encoder/ms-marco-MiniLM-L12-v2 33.4 M 6.1 M
Mono-Decoder hf.co/castorini/monot5-base-msmarco-10k 222.0 M 26.0 M
@@
For the second experiment, we simulate a shallow train set by stratified sampling of $k\in\{64,128,192,256\}$ documents per topic, of which, on average, 12.5% are relevant, depending on availability of relevant documents in the pool. We then finetune topic-specific judge LoRA adapters based on the mono decoder model, using the same hyperparameters as before. We conduct experiments both on the previously used topics of TREC Robust04, and additionally on 29 and 28 topics with more than 50 relevant documents from the TREC Deep Learning 2019 (DL19, [7]) and 2020 (DL20, [6]) passage ranking tracks. Topics with higher counts of relevance labels were chosen to ensure sufficient label variety in both train and test sets even under subsampling conditions.
@@
To compare our proposed judge adapters to established LLM-as-a-judge approaches, we compute relevance labels for the complete TREC DL20 passage pool, using all available topics. We apply two established zero-shot prompts for relevance judgment: (1) UMBRELA [28] (yielding graded relevance labels which we cast to binary), (2) best human-written prompt from Arabzadeh and Clarke [2] (directly yielding binary relevance labels). Furthermore, we test if access to topic-specific information yields improvements in LLM-as-a-judge by adding 8 few-shot examples (4 relevant, 4 non-relevant, in random order) to each prompt. Examples are drawn from the training set of our judge adapter for $t=256$ for comparability. We limit few-shot examples to 8 due to inference cost outweighing additional effectiveness benefit. Inference was carried out using GPT-4o, which Arabzadeh and Clarke [2] showed to be performing best among tested LLMs for relevance judgment.
5 Results & Discussion

### v2

@@ Abstract.
The unjudged document problem, where systems that did not contribute to the original judgement pool may retrieve documents without a relevance judgement, is a key obstacle to the reuseability of test collections in information retrieval. While the de facto standard to deal with the problem is to treat unjudged documents as non-relevant, many alternatives have been proposed, such as the use of large language models (LLMs) as a relevance judge (LLM-as-a-judge). However, this has been criticized, among other things, as circular, since the same LLM can be used as the ranker and the judge. We propose to train topic-specific relevance classifiers instead: By finetuning monoT5 with independent LoRA weight adaptation on the judgments of a single assessor for a single topic’s pool, we align it to that assessor’s notion of relevance for the topic. The system rankings obtained through our classifier’s relevance judgments achieve a Spearmans’ $\rho$ correlation of >0.94>0.94 with ground truth system rankings. As little as 128 initial human judgments per topic suffice to improve the comparability of models, compared to treating unjudged documents as non-relevant, while achieving more reliability than existing LLM-as-a-judge approaches. Topic-specific relevance classifiers are thus a lightweight and straightforward way to tackle the unjudged document problem, while maintaining human judgments as the gold standard for retrieval evaluation. Code, models, and data are made openly available. [^1] https://github.com/webis-de/conf26-topic-adapters-as-judges
1. Introduction
@@
An ad hoc retrieval experiment usually consists of a large collection of documents, several dozen search topics each specifying an information need, and relevance judgments on a representative sample of the documents for each topic (cleverdon1967cranfield; voorhees2018building). For large collections, the manual effort involved in collecting relevance judgments usually prevents a complete assessment of all documents for all topics. Furthermore, most documents in a collection are typically not relevant to a specific topic. To find a compromise between evaluation costs and experimental validity, it is therefore necessary to build a focused subset of the documents for each topic, obtained by combining the top-$k$ documents from the various search engines to be evaluated, i.e., pooling, where $k$ can range from a handful to a few hundred documents per search engine.
Figure 1. Conceptual illustration of how the naïve basline, the LLM-as-a-judge paradigm, and our proposed use of topic-specific judge adapters each operationalize the unjudged document problem.
@@
While this evaluation setup allows for a valid and fair comparison between the systems that initially contribute to the document pool of each topic, any new systems that are implemented after the experiment is complete are at a disadvantage (lipani2016fairness; parry2025variations). These often retrieve documents that are not yet included in any given topic’s pool, leading to the so-called unjudged document problem, namely the still mostly unresolved question of how to deal with such documents. Heuristic solutions such as simply omitting these documents, or deeming them non-relevant, have been proposed and frequently applied. However, none of the heuristics put new and initial systems on an equal footing. As research progresses and more unjudged but potentially relevant documents are uncovered by new retrieval systems, past evaluation experiments and test collections lose the capability to reliably distinguish between them, diminishing their reusability. At the same time, updating a test collection with new judgments is costly and the new judgments for a topic may be incongruent with those from the previous judges (voorhees2022can).
@@
Recent work has proposed to simply prompt instruction-tuned large language models (LLMs) to supplement or completely replace human judgments (abbasiantaeb2025improving; upadhyay2025large; arabzadeh:2025; rahmani2025judging; abbasiantaeb2024can; upadhyay2024umbrela), collectively called LLM-as-a-judge approaches. However, any LLM used as a judge for evaluation (together with its judgment prompt) can simultaneously be used as a ranker in a retrieval system (soboroff2025dont; dietz2025principles). This may lead to an LLM confirming its own ranking decisions (panickssery2024llm; wataoka2024self) during relevance judgment, biasing evaluation results (ye2025justice) in favor of a particular LLM. Moreover, since LLMs inevitably encode social biases picked up from their pre-training data, evaluating retrieval models with LLMs may amplify them in retrieval models (dai2024bias).
@@
We show that the key to address both, the unjudged document problem, and the bias due to using LLMs for both ranking and evaluation, is to train one relevance classifier for each topic based on the judgments of a single human assessor. A topic-specific classifier can only be used to judge documents for that topic, and it ideally closely approximates the notion of relevance originally applied by the human assessor. To demonstrate the effectiveness of this approach, we conduct experiments using the TREC Robust04 (voorhees2004overview), the TREC Deep Learning 2019 (craswell2019overview), and the TREC Deep Learning 2020 (craswell2020overview) passage ranking test collections. We show that the system rankings produced by our topic-specific relevance classifiers strongly correlates with the rankings produced by the available human judgments. We analyze our approach with respect to shallow vs. deep human judgments, and in comparison to zero and few-shot LLM-as-a-judge approaches. Our relevance classifier correlates reliably with human judgments even under limited judgment depth, while outperforming LLM-as-a-judge approaches. These results strongly indicate the feasibility of using topic-specific relevance classifiers to automate retrieval evaluation after bootstrapping one classifier for every topic from gold standard human judgments.
2. Related Work
@@
We review the literature, relating our work to attempts at automating relevance judgments, with a particular focus on recent approaches, and how they differ from ours.
@@
The notion of treating document ranking as a classification problem has been established early in information retrieval with the binary independence model (robertson:1976). Despite the fact that this paradigm to ranking was, for a long time, less effective than other approaches, the idea of exploiting relevance judgments to tune relevance classifiers has also been proposed a long time ago, using logistic regression (gey:1994), investigating active learning as a means to minimize the number of human judgments (lewis:1994), and more recently, using modern re-ranking models (meng2026rerankers). Unlike our approach, this notion of ranking attempts to train a generic relevance classifier for ranking instead of judging, which requires a lot of judgments, since the notion of relevance may differ wildly between topics and between human assessors. Later learning to rank approaches succeeded only once large query logs have been identified as source of implicit relevance feedback (joachims:2002).
Topic-specific relevance classifiers for ranking have been studied at the TREC Filtering tracks 1994 to 2002 (robertson:2002), where participants trained many different kinds of classifiers to filter (classify) documents by their relevance with respect to various different settings. Since each TREC topic is typically assigned to a single assessor, this implicitly means that all these classifiers also approximated the notion of relevance of that specific assessor, just like we propose in this paper.
Efficient Relevance Annotation
Assisting human assessors with automatic means during relevance annotation to increase the efficiency of judgment efforts has been a long-standing problem in information retrieval (soboroff2001ranking). Documents worthwhile to annotate can be identified before labeling commences, e.g., through metric-based heuristics (moffat2006strategic; carterette2006minimal) or LLM-assisted annotation (takehi2025llm). Or, akin to active learning, during labeling in iterative rounds (soboroff2021overview), e.g., by enabling eDiscovery through technology-assisted review (cormack:2014), at the TREC Total Recall tracks 2015 and 2016 (grossman:2016), and by training classifiers to judge documents outside domain-specific tasks (zhang2022survey). We envision the approach described in this paper not as an alternative to such efficient annotation techniques, but complementary to. Our approach allows for more reliable evaluation of test collections after the judgments have been collected, with many ways the initial judgments could be collected.
@@
LLMs have been demonstrated numerous times to have a general sense of relevance (pradeep2023rankzephyr; pradeep2023rankvicuna). Many studies have thus investigated their applicability at collecting relevance judgments (abbasiantaeb2025improving; upadhyay2025large; rahmani2025judgeblender; rahmani2025judging; abbasiantaeb2024can; upadhyay2024umbrela) or even generating partial or fully synthetic test collections (bonifacio2022inpars; rahmani2025syndl; rahmani2024synthetic). The general technique is to develop a prompt that encodes assessment instructions and to input this prompt alongside a query and unjudged documents. We argue that instead, topic specialization (i.e., learning from as much topic expertise as can be obtained) is key to automating relevance judgment in Cranfield-style evaluations, which the current zero-shot approach of LLM-as-a-judge forgoes in favor of generic instructions. While these studies have shown LLM judges to be correlated with human assessment, there are three main issues that current approaches do not address, namely that an LLM-as-a-judge (1) exhibits biases and reliability issues (soboroff2025dont; panickssery2024llm; wataoka2024self; ye2025justice), (2) is highly susceptible to prompt variations (arabzadeh:2025), and (3) can lead to invalid conclusions about evaluation, especially due to circularity (dietz2025principles; soboroff2025dont). Our approach is similar to the idea of LLMs-as-a-judge in that the judgment instructions are encoded into the model. However, rather than the prompt being used as the medium for storing what should be relevant, we directly encode what should be relevant into the parameters of the judging model. In doing so, we avoid the common pitfalls associated with LLMs-as-a-judge.
@@
The idea to predict relevance judgments and expand the overall set of relevance judgments was first proposed by (buettcher2007reliable). The general idea was to train a binary SVM classifier using TF–IDF features from documents to determine relevance. There are three main issues with these experiments that we address in this paper: (1) Their approach did not work well for deep assessments. Our approach effectively predicts relevance at much deeper depth. (2) Their approach relies heavily on lexical features. Today, effective ranking models use much more complex, semantic scoring functions and our approach models semantics much more effectively. (3) Their approach trains classifiers from scratch, while we opt for adaptation of existing ranking models, leveraging their strong relevance priors.
3. Topic-Specific Relevance Classifiers for Judging
@@
We leverage pretrained ranking models as initialization rather than training classifiers from scratch. Models pretrained on large-scale relevance data provide strong semantic representations and implicit priors over document-query relevance (pradeep2023rankzephyr; pradeep2023rankvicuna). We experiment with base ranking models from three archetypes: bi-encoders (representation-based), cross-encoders (interaction-based), and mono-decoders (generative scoring). Yet, full fine-tuning per topic is computationally prohibitive and produces large model weights unfit for distribution. We thus rely on LoRA (hu2022lora) to inject topic-specific behavior while keeping base model weights frozen. LoRA learns a low-rank approximation of the weight modifications attained during training. This allows each topic to receive independent adapter weights with no parameter sharing, preventing cross-topic interference. It further allows for lightweight distribution: one base model, many shareable topic-specific adapters.
4. Experimental Setup
We conduct three experiments. The first investigates if, given a deep pool of judged data, training judge models is feasible. The second investigates how this changes if the pool changes to a shallow one, and how the predicted relevance judgments produced by a finetuned LoRA adapter (“judge adapter”) impact ranking evaluation results. Finally, the third investigates how the approach compares to LLM-as-a-judge methods, and if those can likewise benefit from topic-specific alignment through few-shot prompting.
@@ 4. Experimental Setup
We select 47 topics with at least 100 relevant documents from the deeply judged official eval set of the TREC Robust04 dataset (voorhees2004overview). Annotated graded relevance labels are cast to binary relevance. This yields, on average, $1350.31(444.23)$ non-relevant and $183.02(78.76)$ relevant documents per topic. We divide the available judgments in each topic into a train and a test set with a 80/20 split, with stratified sampling per label to ensure similar relevancy distribution in both sets. Then, for each of the three pointwise model archetypes bi-encoder (reimers2019sentence), cross-encoder (reimers2019sentence), and mono-decoder (nogueira2020document), we select representative pretrained models as listed in Table˜1. We separately finetune topic-specific judge models for each topic and archetype, over 10 epochs, with a batch size of 64 query/doc pairs of a combined sequence length of 512 tokens, and a learning rate of 1e-4. We train two version per topic: a native finetuning of the full model weights, and a LoRA adapter (hu2022lora) with parameters $r=64$ and $\alpha=128$. The number of trainable parameters in each setting and for each model is also listed in Table˜1. Given the high class imbalance in each topic, we use a weighted MSE loss, with weights 0.95 and 0.05 for relevant/non-relevant labels. This corresponds to the inverse average relevance distribution in judgments.
Table 1. Overview of the models investigated as topic-specific judges in this paper.
@@
Bi-Encoder all-MiniLM-L6-v2 22.7 M 5.4 M
Cross-Encoder ms-marco-MiniLM-L12-v2 33.4 M 6.1 M
Mono-Decoder monot5-base-msmarco-10k 222.0 M 26.0 M
Table 2. Overview of the LLMs backbones investigated for LLM-as-a-judge in this paper.
Model Checkpoint Param. Reas.
Ministral-3 Ministral-3-14B-Instruct-2512 114 B
Qwen-3 Next Qwen3-Next-80B-A3B-Instruct 180 B
GPT-OSS gpt-oss-120b 120 B
MiniMax-M2.1 MiniMax-M2.1 229 B
GPT-4o gpt-4o-2024-11-20
@@
For the second experiment, we simulate a shallow train set by stratified sampling of $k\in\{64,128,192,256\}$ documents per topic, of which, on average, 12.5% are relevant, depending on availability of relevant documents in the pool. We then finetune topic-specific judge LoRA adapters based on the mono decoder model, using the same hyperparameters as before. We conduct experiments both on the previously used topics of TREC Robust04, and additionally on 29 and 28 topics with more than 50 relevant documents from the TREC Deep Learning 2019 (DL19, (craswell2019overview)) and 2020 (DL20, (craswell2020overview)) passage ranking tracks. Topics with higher counts of relevance labels were chosen to ensure sufficient label variety in both train and test sets even under subsampling conditions.
@@
To compare our proposed judge adapters to established LLM-as-a-judge approaches, we compute relevance labels for the complete TREC DL20 passage pool, using all available topics. We apply two established zero-shot prompts for relevance judgment: (1) UMBRELA (upadhyay2024umbrela) (yielding graded relevance labels which we cast to binary) and (2) the best human-written prompt from (arabzadeh:2025) (directly yielding binary relevance labels). Furthermore, we test if access to topic-specific information yields improvements in LLM-as-a-judge by adding 8 few-shot examples (4 relevant, 4 non-relevant, in random order) to each prompt. Examples are drawn from the training set of our judge adapter for $t=256$ for comparability. We limit few-shot examples to 8 due to inference cost outweighing additional effectiveness benefit. Inference was carried out using varied open LLMs including both reasoning and non-reasoning models (see Table˜2). Additionally, we include the closed model GPT-4o for reference, since (arabzadeh:2025) found it to be performing best for judging tasks.
Figure 2. Effectiveness of our topic-specific ranker approach in terms of F1, precision, recall, and accuracy when fine-tuning versus when using LoRA between different model architectures.
5. Results & Discussion